OpenStudy (anonymous):

Factor completely: 18x^2 - 21x - 15

OpenStudy (anonymous):

First, you need to factor out a 3

OpenStudy (anonymous):

That leaves you with 3(6x^2-7x-5)

OpenStudy (anonymous):

Then continue factoring 6x^2-7x-5 which gives you 3(2x+1)(3x-5)

OpenStudy (anonymous):

That should be your answer

OpenStudy (ybarrap):

Starting with \(\tt 18x^2-21x-15\) Bring out your 3: $$\tt =3(6x^2-7x-5)$$ Find factors that add to -7 but multiply to give you 30 (i.e 6*5): $$\tt =3\frac{(6x + 3)(6x - 10)}{6}$$ Bring out common terms: $$\tt =3\frac{(3(2x + 1)2(3x - 5)}{6}$$ Reduce $$\tt =3(2x + 1)(3x - 5)$$ Done

OpenStudy (ybarrap):

*give you -30